OpenStudy (anonymous):

5 over 13 times T = –9?

OpenStudy (anonymous):

I am assuming \[(5/13)x=-9\] 5x=-9*13=-117 X=-117/5

OpenStudy (anonymous):

Check: 5/13*(-117/5)=-9 True